One mole of diatomic gas is heated at 1atm from 0 k to 100 k .if the gas constant R=8.32joule/mole K,change in internal energy is approximately 

Grade:11

2 Answers

Arun
5 years ago
Internal energy = f/2 × nR∆T

for di atomic gas f= 5

given. 

n =1 

and 

∆T = 100-0 = 100

so 

f = 5 gh/2 ×1× 25/3 × 100

= 2083.3J
Rajdeep
5 years ago
HELLO THERE!
 
There’s a formula:
 
\triangle U = nC_{v}\triangle T
 
Where \small C_{v} is Specific heat at constant volume, and for diatomic gas, it is 
\small C_{v} = \frac{5}{2}R
 
Change in temperature = 100 – 0 = 100K
Number of moles (n) = 1
 
So, we have:
 
\small \triangle U = n C_{v}\triangle T \\\\= 1 \times \frac{5}{2}\times 8.32\times 100 \\\\= 2080 J
 
THANKS!